OpenStudy (anonymous):

evaluate the function f(x)=8x+2 use for x f(x+3)

OpenStudy (anonymous):

i think the awnser is 8x+2 not sure

OpenStudy (radar):

f(x)=8x+2 f(x+3)=8(x+3) + 2=8x+24+2 = 8x+26

OpenStudy (anonymous):

thanks

OpenStudy (radar):

You are welcome.